# Modeling Spatio-Temporal Transport: From Rigid Advection to Realistic Dynamics

### v3

@@ 4.1 Framework
For a fixed wavenumber $\mathbf{k}$ and velocity $\mathbf{v}$, the spectrum of a FF (4) consists of the spectrum $S_{XX}(\mathbf{k})$ of the spatial random field multiplied by a Dirac delta centered in $\omega=-\mathbf{k}^{T}\mathbf{v}$. Such separable structure characterizing $S_{ZZ}(\mathbf{k},\omega)$ makes it particularly convenient for us to modify the part of the spectrum that depends on velocity $\mathbf{v}$. For instance, replacing the Dirac deltas with a “milder” singularity would ensure that the spectrum includes more frequencies neighbouring $\omega=-\mathbf{k}^{T}\mathbf{v}$. This introduces temporal decorrelation into the model, which gains the flexibility to represent more complex and realistic phenomena where features naturally change, dissipate, or evolve over time.
@@
Since $|e^{i\theta}|=1$ for any real $\theta\in\mathbb{R}$, the inequality simplifies to $|c_{ZZ}(\mathbf{v}\tau,\tau)|\leq c_{ZZ}(0,0)$. Equality can only hold if the argument of the integral, $S_{ZZ}(\mathbf{k},\omega)e^{2\pi i\tau(\mathbf{k}^{\top}\mathbf{v}+\omega)}$, has a constant phase for all $(\mathbf{k},\omega)$ where the spectrum $S_{ZZ}$ is non-zero. Given that $S_{ZZ}(\mathbf{k},\omega)$ is a real and non-negative function, it does not affect the phase. Therefore, the phase of the complex exponential term $e^{2\pi i\tau(\mathbf{k}^{\top}\mathbf{v}+\omega)}$ must be constant, which implies that its exponent, $2\pi\tau(\mathbf{k}^{\top}\mathbf{v}+\omega)$, must be constant. This condition must hold for all $(\mathbf{k},\omega)$ with non-zero spectral energy. If we consider the point $(\mathbf{k}=\mathbf{0},\omega=0)$, which must also satisfy this relation, we find that $2\pi\tau(0+0)=0$. Thus, with the constant being zero, the condition for equality becomes $\mathbf{k}^{\top}\mathbf{v}+\omega=0$ on the set of wavenumber-frequency $\{(\mathbf{k},\omega)\in\mathbb{R}^{2}\times\mathbb{R}:S_{ZZ}(\mathbf{k},\omega)>0\}$, implying $\omega=-\mathbf{k}^{\top}\mathbf{v}$ in the same set. This means the spectrum $S_{ZZ}$ must be entirely supported on the hyperplane defined by $\omega=-\mathbf{k}^{\top}\mathbf{v}$, which is precisely the spectral definition of a FF.
@@
A Damped Frozen Field is a zero-mean Gaussian process characterized by a persistence parameter $\alpha>0$ and a damping parameter $\beta\geq 0$, whose spectrum is defined as
@@
This model captures the interplay between advection and physical decorrelation through its two key parameters. The persistence parameter, $\alpha$, controls the strength of the process’s memory along the advective path. The damping parameter, $\beta$, prevents the singularity of a FF, and governs the geometric shape of the spectral decay around the advection line. Notice that persistence parameter $\alpha$ must be positive. This ensures that the model describes a system where energy decays at higher frequencies, which is a necessary condition for a well-behaved physical process. Furthermore, for the total variance of the process to be finite, i.e. $c_{ZZ}(0,0)=\iiint S_{ZZ}(\mathbf{k},\omega)\,\mathrm{d}\omega\,\mathrm{d}^{2}\mathbf{k}<\infty$, the spectrum must be integrable. This typically imposes a stronger lower bound on $\alpha$, depending on the behavior of the spatial spectrum $S_{XX}(\mathbf{k})$ as $\mathbf{k}\to 0$. We will discuss further the roles of parameters $\alpha$ and $\beta$ in Section 4.2.
@@ 4.2 Role of the Parameters
The spectrum in Equation (19) is shaped by the term $B(\mathbf{k},\omega;\mathbf{v},\beta)^{-\alpha}$. The power-law form of this term is characteristic of a long-memory process, but with the critical distinction that its behavior is centered on the advection line $\omega=-\mathbf{k}^{\top}\mathbf{v}$ rather than the origin. This formulation defines a space-time long-memory process with advection, whose characteristics are non-trivial due to the non-separable nature of $B(\mathbf{k},\omega;\mathbf{v},\beta)$. This term acts as a filter that intrinsically links a wave’s temporal evolution to its spatial wavenumber, which is a physically realistic assumption. The filter reshapes the initial spatial spectrum by amplifying the signal where $B(\mathbf{k},\omega;\mathbf{v},\beta)<1$, and attenuating it where $B(\mathbf{k},\omega;\mathbf{v},\beta)>1$. The geometry of this amplification region is primarily controlled by the damping parameter $\beta$, while the persistence parameter $\alpha$ dictates the intensity of the effect.
@@
Figure 5: Heatmap of the logarithm of $B(\mathbf{k},\omega;\mathbf{v},\beta)^{-\alpha}$ for different choices of $\alpha$ (rows) and $\beta$ (columns). Level curves are shown in a gradient of pinks, where lightest correspond to highest values. The dashed black line correspond to $\omega=-\mathbf{k}^{\top}\mathbf{v}$.
@@ 5 Conclusions
Our first strategy led to the development of the Distributed FF and the Evolving FF models. The Distributed FF, by representing the process as a superposition of fields with different advective velocities, provides a natural framework for modeling phenomena driven by multiple transport influences. We demonstrated that this model is not only capable of producing complex dynamics, such as those of a tropical cyclone, but also that it provides a direct stochastic analogue to the advection-diffusion PDE. The Evolving FF, in contrast, allows the velocity to vary locally in space and time, offering a powerful tool for simulating processes within complex, non-uniform flow fields. For both models, we derived key properties, including their propagation paths, and provided efficient simulation algorithms based on modifications of the FF simulation.
@@
The frameworks established in this paper open several promising directions for future investigation. First, while our work has focused on model definition and simulation, an avenue for future research is the development of robust statistical methods for fitting these models to observational data. This includes estimating the velocity distribution for a Distributed FF, the latent velocity field for an Evolving FF, or the spectral parameters for a Damped FF. Techniques ranging from Whittle likelihood estimation (guillaumin2019debiased) to hierarchical Bayesian modeling (see for instance rue2007 and chagneau2011) could be explored for this purpose. Once estimation procedures are developed, a crucial line of inquiry will be the direct comparison of these models against competing approaches in the literature. Such a comparison should focus on predictive performance and forecasting capabilities, which are critical for real-world applications in fields like climatology, atmospheric science, and oceanography. Second, many environmental systems involve multiple, co-evolving variables, like temperature and pressure, and pollutants and wind velocity. Extending our models to a multivariate setting, where several fields are jointly advected and interact, is a critical next step. The multivariate FF model of salvana2022jasa provides a foundational starting point for developing multivariate versions of the Distributed, Evolving, and Damped FF models. Finally, the Damped FF model introduces a form of long memory centered on the advection path. A deeper theoretical investigation could formally connect this model to the theories of fractional Brownian motion (e.g., flandrin1989; qian1998two; stein2002fast; lilly2017fractional; didier2018domain) and spatial long-memory processes (robinson2020spatial). For instance, it would be valuable to understand the interaction between the long-memory characteristics of the initial spatial field $X_{S}(\mathbf{s})$ and the long-memory properties induced by the advective spectral filter.
@@ Appendix A Algorithms
This section outlines the simulation algorithms for the models presented in the main paper. Algorithm 3 shows how to simulate a Distributed Frozen Field as a convex combination of $n_{v}$ simulated Frozen Fields; each component $i=1,\dots,n_{v}$ is advected with a unique velocity $\mathbf{v}_{i}=\mathbf{M}_{i}\mathbf{v}$ and assigned a weight $p_{i}=p(\mathbf{M}_{i}\mathbf{v})$. For the Evolving Frozen Field, Algorithm 4 adapts the simulation procedure of a Frozen Field by substituting the constant advection velocity with a vector field $\mathbf{v}(\mathbf{s},t)$ that varies for each spatial location $\mathbf{s}$ and time point $t$. Finally, Algorithm 5 presents the spectral simulation of a Damped Frozen Field, which is generated from a spectrum constructed using the spatial field’s spectrum $S_{XX}$, and a filter whose shape is determined by a persistence parameter $\alpha>0$ and a damping parameter $\beta\geq 0$.
@@
2:Spatial spectral density $S_{XX}$, persistence parameter $\alpha>0$, damping parameter $\beta\geq 0$.

### v4

@@ 4.1 Framework
For a fixed wavenumber $\mathbf{k}$ and velocity $\mathbf{v}$, the spectrum of a FF (4) consists of the spectrum $S_{XX}(\mathbf{k})$ of the spatial random field multiplied by a Dirac delta centered in $\omega=-\mathbf{k}^{\top}\mathbf{v}$. Such multiplicative structure characterizing $S_{ZZ}(\mathbf{k},\omega)$ makes it particularly convenient for us to modify the part of the spectrum that depends on velocity $\mathbf{v}$. For instance, replacing the Dirac deltas with a “milder” singularity would ensure that the spectrum includes more frequencies neighbouring $\omega=-\mathbf{k}^{\top}\mathbf{v}$. This introduces temporal decorrelation into the model, which gains the flexibility to represent more complex and realistic phenomena where features naturally change, dissipate, or evolve over time.
@@
Since $|e^{i\theta}|=1$ for any real $\theta\in\mathbb{R}$, the inequality simplifies to $|c_{ZZ}(\mathbf{v}\tau,\tau)|\leq c_{ZZ}(0,0)$. Equality can only hold if the argument of the integral, $S_{ZZ}(\mathbf{k},\omega)e^{2\pi i\tau(\mathbf{k}^{\top}\mathbf{v}+\omega)}$, has a constant phase for all $(\mathbf{k},\omega)$ where the spectrum $S_{ZZ}$ is non-zero. Given that $S_{ZZ}(\mathbf{k},\omega)$ is a real and non-negative function, it does not affect the phase. Therefore, the phase of the complex exponential term $e^{2\pi i\tau(\mathbf{k}^{\top}\mathbf{v}+\omega)}$ must be constant, which implies that its exponent, $2\pi\tau(\mathbf{k}^{\top}\mathbf{v}+\omega)$, must be constant. This condition must hold for all $(\mathbf{k},\omega)$ with non-zero spectral energy. If we consider the point $(\mathbf{k}=\mathbf{0},\omega=0)$, which must also satisfy this relation, we find that $2\pi\tau(0+0)=0$. Thus, with the constant being zero, the condition for equality becomes $\mathbf{k}^{\top}\mathbf{v}+\omega=0$ on the set of wavenumber-frequency $\{(\mathbf{k},\omega)\in\mathbb{R}^{2}\times\mathbb{R}:S_{ZZ}(\mathbf{k},\omega)>0\}$, implying $\omega=-\mathbf{k}^{\top}\mathbf{v}$ in the same set. This means the spectrum $S_{ZZ}$ must be entirely supported on the hyperplane defined by $\omega=-\mathbf{k}^{\top}\mathbf{v}$, which is precisely the spectral definition of a FF, for which $\rho(\mathbf{v}\tau,\tau)=1$.
@@
A Damped Frozen Field is a zero-mean Gaussian process characterized by a persistence parameter $\alpha>1/2$ and a damping parameter $\beta\geq 0$, whose spectrum is defined as
@@
This model captures the interplay between advection and physical decorrelation through its two key parameters. The persistence parameter, $\alpha$, controls the strength of the process’s memory along the advective path. The condition $\alpha>1/2$ ensures that the spectral density has finite integral over frequency. This is because, as $|\omega|\to\infty$, we have $B(\mathbf{k},\omega;\mathbf{v},\beta)\sim\omega^{2}(1+\beta^{2})$, so that $S_{ZZ}(\mathbf{k},\omega)\sim|\omega|^{-2\alpha}$. The integral $\int_{\mathbb{R}}|\omega|^{-2\alpha}d\omega$ converges if and only if $2\alpha>1$, yielding the requirement $\alpha>1/2$. Under such parameter restriction, the model describes a system where energy decays at higher frequencies, which is a necessary condition for a well-behaved physical process. The damping parameter $\beta\geq 0$ prevents the singularity of a FF at zero frequency, and governs the geometric shape of the spectral decay around the advection line. Furthermore, for the total variance $c_{ZZ}(0,0)=\iiint S_{ZZ}(\mathbf{k},\omega)\,\mathrm{d}\omega\,\mathrm{d}^{2}\mathbf{k}$ to be finite, the spatial spectrum $S_{XX}(\mathbf{k})$ must also be integrable over wavenumber $\mathbf{k}$, which is a standard assumption for well-defined spatial random fields.
The following result shows that the spectrum of the Damped FF corresponds to a valid covariance function.
@@
This spectral construction admits a rigorous interpretation within the theory of SPDE-based stationary random fields proposed by carrizo2022. In their framework, the spatio-temporal spectral density is derived from the symbol function $g(\mathbf{k},\omega)$ of a differential operator. Our damping term $B(\mathbf{k},\omega;\mathbf{v},\beta)^{-\alpha}$ plays an analogous role to the inverse of the squared symbol function $|g(\mathbf{k},\omega)|^{-2}$ in their Theorem 1. Consequently, the Damped FF can be viewed as the stationary solution to a fractional evolution SPDE driven by a source term with spatial spectrum $S_{XX}$. This connects our spectral approach to the advection-diffusion SPDEs studied by carlotto2024. However, while carlotto2024 rely on Finite Element discretizations to handle spatially varying parameters and scattered data, our spectral formulation leverages the Fast Fourier Transform for computationally efficient simulation on regular grids.
@@ 4.2 Role of the Parameters
The spectrum in Equation (20) is shaped by the term $B(\mathbf{k},\omega;\mathbf{v},\beta)^{-\alpha}$. The power-law form of this term determines the smoothness and decay of temporal correlations, centered on the advection line $\omega=-\mathbf{k}^{\top}\mathbf{v}$. While this does not constitute a long-memory process in the strict sense of a spectral divergence (robinson2020spatial), this formulation allows for a tunable degree of temporal persistence. In particular, the term $B(\mathbf{k},\omega;\mathbf{v},\beta)$ acts as a filter that intrinsically links a wave’s temporal evolution to its spatial wavenumber, which is a physically realistic assumption. The filter reshapes the initial spatial spectrum by amplifying the signal where $B(\mathbf{k},\omega;\mathbf{v},\beta)<1$, and attenuating it where $B(\mathbf{k},\omega;\mathbf{v},\beta)>1$. The geometry of this amplification region is primarily controlled by the damping parameter $\beta$, while the persistence parameter $\alpha$ dictates the intensity of the effect.
@@
Figure 7: Heatmap of the logarithm of $B(\mathbf{k},\omega;\mathbf{v},\beta)^{-\alpha}$ for different choices of $\alpha$ (rows) and $\beta$ (columns). Level curves are shown in a gradient of pinks, where lightest correspond to highest values. The dashed black line correspond to $\omega=-\mathbf{k}^{\top}\mathbf{v}$. Note that $\alpha=0.5$ represents the theoretical limit of the model’s validity, where the spectral density is at the threshold of integrability.
@@ 5 Conclusions
Our first strategy led to the development of the Distributed FF and the Evolving FF models. The Distributed FF, by representing the process as a superposition of fields with different advective velocities, provides a natural framework for modeling phenomena driven by multiple transport influences. We demonstrated that this model is not only capable of producing complex dynamics, such as those of a tropical cyclone, but also that it provides a direct stochastic analogue to the advection-diffusion PDE. The Evolving FF, in contrast, allows the velocity to vary locally in space and time, offering a powerful tool for simulating processes within complex, non-uniform flow fields, like the evolving rainbands during Hurricane Florence. For both models, we derived key properties, including their propagation paths, and provided efficient simulation algorithms based on modifications of the FF simulation.
@@
The frameworks established in this paper open several promising directions for future investigation. While this work has focused on model definition and simulation, an important avenue for future research is the development of robust statistical methods for fitting these models to observational data. Each of the three proposed models naturally suggests specific estimation strategies tailored to its mathematical structure. For the Distributed FF, the explicit analytical form of the covariance function in Equation (7) enables classical spatial statistics approaches. Minimum contrast estimation (Cressie2011) could be employed by matching the theoretical covariance to an empirical estimate, with parameters governing the velocity distribution optimized to minimize a distance criterion. Alternatively, composite likelihood methods (varin2011) offer a computationally tractable approach for large datasets by constructing a pseudo-likelihood from pairwise or block contributions, circumventing the need to evaluate the full joint density. For the Evolving FF, the central inferential challenge is recovering the latent velocity field $\mathbf{v}(\mathbf{s},t)$ from observed realizations of the process. This problem has a direct analogue in computer vision, where optical flow methods (see for instance horn1981) estimate motion fields by minimizing discrepancies between consecutive frames subject to smoothness constraints. This approach was used in the application to Hurricane Florence data. Within a parametric framework, another strategy is to represent the velocity field using a tensor product of known basis functions, specifically spatial and temporal basis functions. Then, the basis coefficients become the target of the estimation. This reduces the infinite-dimensional problem to a finite-dimensional one, amenable to penalized least squares or Bayesian inference with appropriate priors on the coefficients (wood2017generalized; ramsay2005functional). Moreover, physical constraints, such as incompressibility ($\nabla\cdot\mathbf{v}=0$), can be incorporated through the choice of penalty terms (corpetti2002). Finally, the Damped FF is particularly amenable to frequency-domain inference due to its explicit spectral density in Equation (20). Whittle likelihood estimation (whittle1954) would provide a natural framework, approximating the log-likelihood as a function of the periodogram and the theoretical spectrum. This approach is computationally attractive because both the periodogram and the parametric spectrum can be evaluated efficiently via the Fast Fourier Transform, making it feasible to estimate the persistence parameter $\alpha$, the damping parameter $\beta$, and the advection velocity $\mathbf{v}$ even for large spatio-temporal datasets. Recent advances in debiased Whittle estimation (guillaumin2019debiased) address edge effects and improve finite-sample performance, which would be particularly relevant given the bounded domains typical in applications. Second, many environmental systems involve multiple, co-evolving variables, like temperature and pressure, and pollutants and wind velocity. Extending our models to a multivariate setting, where several fields are jointly advected and interact, is a critical next step. The multivariate FF model of salvana2022jasa provides a foundational starting point for developing multivariate versions of the Distributed, Evolving, and Damped FF models. Finally, the Damped FF model introduces a structure of tunable persistence centered on the advection path. A deeper theoretical investigation could formally connect this model to the theories of fractional dynamics and regularity found in the fractional Brownian motion literature (e.g., flandrin1989; qian1998two; stein2002fast; lilly2017fractional; didier2018domain). For instance, it would be valuable to understand the interaction between the fractional characteristics of the initial spatial field $X_{S}(\mathbf{s})$ and the regularity properties induced by the advective spectral filter.
@@ Appendix A Algorithms
This section outlines the simulation algorithms for the models presented in the main paper. Algorithm 3 shows how to simulate a Distributed Frozen Field as a convex combination of $n_{v}$ simulated Frozen Fields; each component $i=1,\dots,n_{v}$ is advected with a unique velocity $\mathbf{v}_{i}=\mathbf{M}_{i}\mathbf{v}$ and assigned a weight $p_{i}=p(\mathbf{M}_{i}\mathbf{v})$. For the Evolving Frozen Field, Algorithm 4 adapts the simulation procedure of a Frozen Field by substituting the constant advection velocity with a vector field $\mathbf{v}(\mathbf{s},t)$ that varies for each spatial location $\mathbf{s}$ and time point $t$. Finally, Algorithm 5 presents the spectral simulation of a Damped Frozen Field, which is generated from a spectrum constructed using the spatial field’s spectrum $S_{XX}$, and a filter whose shape is determined by a persistence parameter $\alpha>1/2$ and a damping parameter $\beta\geq 0$.
@@
2:Spatial spectral density $S_{XX}$, persistence parameter $\alpha>1/2$, damping parameter $\beta\geq 0$.